If I have a 20 qt radiator containing a 80% antifreeze solution. How much of the solution should I drain and replace with pure water to get a solution that is 50% antifreeze?
Answer by mananth(16946)

Let 82% solution required be x
percent ---------------- quantity
82 ---------------- x
water 0% ---------------- 20-x
50.00% ---------------- 20
...
82x+0(20 -x)=50*20
82x=1000
/ 82
x=12.2 qt. of82.00%
Balance 7.8 qt of water.
7.8 qt. of 82% will have to be drained and replaced with water